One bus left a terminal traveling east at 80 kilometers per hour. A second bus left an hour and a half later traveling west at ninety kilometers per hour. At what time will the buses be 460 kilometers apart?

Please help, thanks.

Answer by jim_thompson5910(35256)

Let t = time

First bus:

d = rt

d = 80t

-------------------------------------------------------

Second Bus:

d = rt

d = 90(t-1.5)

Add up the two distances that the buses travel, set that equal to 460, then solve for t to get

80t + 90(t - 1.5) = 460

80t + 90t - 135 = 460

170t - 135 = 460

170t = 460 + 135

170t = 595

t = 595/170

t = 3.5

So it will take 3.5 hours (3 and a half hours, or, 3 hours and 30 minutes)

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------